# Reconstructing a Heavy Neutral Lepton at the LHC

### v3

@@ 4 Discussion
In any search for a new particle it is critical to find the optimal kinematical variable that discriminates between signal and background. Here we have discussed how to reconstruct the mass of a HNL in the trilepton plus $p_{\rm T}^{\rm miss}$ channel. This possibility may have been overlooked in several CMS analyses [15, 16], where they use a combination of observables optimized with machine learning techniques that distributes signal and background in 20 bins. Obviously, when the signal events define a mass peak its search will always be the better choice.
@@
Our estimate does not carry the degree of sophistication in these full experimental analyses, that include next to leading order corrections, the background of photons and heavy mesons misidentified as leptons, or a precise account of the detectors response. However, it seems clear that the appearance of a mass peak should simplify the search and make the details of the calculation less determinant. In summary, although it is a possibility not explored at the LHC yet, the analysis presented here suggests that the search for peaks associated to HNLs in $pp$ collisions could probe mixings $|V_{\ell N}|^{2}$ that are not excluded by PMNS unitarity bounds.

### v4

@@ 4 Discussion
[…]
Here we have discussed the reconstruction of $m_{N}$ in the trilepton plus neutrino channel and how to solve the two-fold degeneracy in the quadratic equation that provides the longitudinal momentum of the neutrino. We find that this equation has no solution in a small fraction (around 10%) of signal events. This is mainly due to the smearing effect introduced by the detector, and also to the different kinematics in signal events where the third lepton comes from a tau lepton. We find interesting that the background includes a much larger fraction (around 25%) of events with no solution to the equation, since this could be used as an extra cut also when studying other observables.
@@
In summary, the mass peak discussed here could complement and simplify the search for an HNL in $pp$ collisions and make the details of the calculation less determinant: even if it is not as clean as the ones in Higgs decays, the peak is largely insensitive to normalization uncertainties related to luminosity, overall cross sections, or selection efficiencies. Although it is a possibility not explored at the LHC yet, our results suggest that its search could help to set collider limits on $|V_{\ell N}|^{2}$ competitive with PMNS unitarity bounds.